i ride 2 hours on a bike. i ride out at 18 mph and ride back at 25 mph
Answer by ewatrrr(24785)

 
Hi
Distance the same at both rates: Let x represent time taken traveling back.
25mphx = 18mph(2-x)
43x = 36
x = 36/43 hr, time traveling back
D = 25*(36/43) = 20.93mi
20.93mi = 18*(2-36/43) = 18(86/43-36/43) = 18*(50/43) =